OpenStudy (anonymous):

The atmospheric pressure 'p' in Pascals, is related to the number of kilometers, 'h' above sea level by the function p(h)=760(2.7)^-0.145h. Find the atmospheric pressure at the height of 10 kilometers. Round answer to nearest tenth. Show all work.

OpenStudy (stormfire1):

This is a simple plug and chug problem: \[p(h)=760*2.5^{-0.145h}\]\[p(h)=760*2.5^{-0.145(10km)}\] Plug that into your calculator to get the answer.